OpenStudy (anonymous):

A rectangular area of 3750 square feet is to be fenced off. Two oppositesides will have fencing that costs $2 per foot and the remaining sides will use fencing that costs $3 per foot. Find the dimensions of the rectangle that will minimize the cost.

OpenStudy (anonymous):

find the factors of 3750 find the perimeter

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

50 x 75 = 3750

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

50 + 50 + 75 + 75 = Perimeter

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait hold on i think i may have messed up lol

OpenStudy (anonymous):

@electrokid can you see if i did this right?

OpenStudy (anonymous):

@electrokid so far i have been able to do it 3 ways but each time i got a different answer

OpenStudy (anonymous):

let the 2$ side be "l" and 3$ side be w \[A=l\times w=3750\\ C=2(2l)+3(2w)\\ C(w)=4\times{3750\over w}+6w \]

OpenStudy (anonymous):

now, maximize C(w) so, its detrivative should be 0

OpenStudy (anonymous):

\[C'(w)=-4\times{3750\over w^2}+6=0\\ 6w^2=15000\\ w^2=2500\\ \boxed{w=50}\\ l={3750\over50}\implies\boxed{l=50} \]

OpenStudy (radar):

You know which dimension you are going to use the hi-dollar fence the 50 ft, the 75 ft dimension will have the 2$ per ft. fencing.